butterbear:

Which is equivalent to 216 1/3? 3 6 36 72

dude:

\(216\frac13\)? Seems like the equation was copied incorrectly, could you check or tell me if a number is formatted a certain way?

butterbear:

yes thats it

butterbear:

its above the right corner on top of the 216 here

butterbear:

563blackghost:

So it is \(\large\bf{216^{\frac{1}{3}}}\)

563blackghost:

The representation of a fraction being the exponent means you'll be finding the root of the number. `216` is being raised to the power of `1/3` you would follow by: \(\large\bf{a^{\frac{m}{n}} = \sqrt[n]{a^{m}}}\) So it would be: \(\large\bf{216^{\frac{1}{3}} ~~~\rightarrow ~~~ \sqrt[3]{216}}\) You plug into a calculator to find the cube root of 216.

butterbear:

\(\color{blue}{yep}\) ]

butterbear:

how do you do the 3

butterbear:

\(\color{red}{its~so~WEIRD}\)

563blackghost:

Following the formula you get 3. \(\LARGE\bf{a^{\frac{\color{green}{m}}{\color{red}{n} }}= \sqrt[\color{red}{n}]{a^{\color{green}{m}}}}\) You follow this. Plug. \(\Large\bf{216^{\frac{\color{green}{1}}{\color{red}{3}}}~~\rightarrow ~~\sqrt[\color{red}{3}]{216^{\color{green}{1}}}~~\rightarrow~~\sqrt[\color{red}{3}]{216}}\)

butterbear:

kk thanks

563blackghost:

no problem